OpenStudy (anonymous):

Show how tan(x+45 degrees) = 6tanx can be written in the form 6tan^2 x-5x+1=0

OpenStudy (anonymous):

tan(x + 45) = 6 tan x \[\frac{ \tan x + \tan 45 }{ 1-\tan x.\tan45 }= 6 tan x\] \[\frac{ \tan x + 1 }{ 1 - \tan x }=6 tan x\] tan x + 1 = 6 tan x(1 - tan x) tan x + 1 = 6 tan x - 6 tan^2 x 6 tan^2 x - 6 tan x + tan x + 1 = 0 6 tan^2 x - 5 tan x + 1 = 0

OpenStudy (anonymous):

Thanks. :)